# Impact of Neutrino Flavor Conversions on Neutron Star Merger Dynamics, Ejecta, Nucleosynthesis, and Multi-Messenger Signals

### v1

@@ III.3 $Y_{e}-\theta$ profiles
In Fig. 9 we show the average electron fractions for the ejecta as a function of the latitude $\theta$, for both the DD2 and SFHo runs. For the DD2 models, the ejecta average $Y_{e}$ increases almost monotonically with latitude. While for the SFHo runs, the average $Y_{e}$ of ejecta shows a small bump around $\theta\approx 20^{\circ}$, which is likely associated with shock heating. The main differences between the no-mixing and mixing runs with the DD2 EoS occur near the equatorial plane and at intermediate latitudes ($\theta<45^{\circ}$), with their average $Y_{e}$ differing by up to $0.1$. In the SFHo runs, the contrast between mixing and no-mixing models is likewise most pronounced in the near-equatorial ejecta, with little variation in the polar regions. The only exception is the SFHo-FFI model, which exhibits up to $\sim 0.05$ lower average $Y_{e}$ than the SFHo-NM model at higher latitudes ($\theta>45^{\circ}$). The differences in the $Y_{e}-\theta$ profiles between mixing and no-mixing models suggest that neutrino flavor conversions have larger impacts on ejecta near the equatorial plane and at intermediate latitudes, than ejecta in the polar regions.
@@ IV Conclusion
In agreement with earlier work Qiu et al. (2025); Li and Siegel (2021); Just et al. (2022), we find that mixing scenarios drive the matter toward more neutron-rich conditions compared to no-mixing models. The neutrino flavor transformations of electron (anti)neutrinos into heavy-lepton types in the inner disk regions boost the total neutrino luminosities at larger radii. While species-dependent hierarchies vary across the EoS and collapse outcomes, we generally find heavy-lepton luminosities amplified in the mixing models relative to the no-mixing case, and vice versa for the electron (anti)neutrinos. Flavor conversions raise the mean energies of electron (anti)neutrinos and lower those of heavy-lepton neutrinos, effectively shifting the decoupling surface of electron-flavor neutrinos inward and that of heavy-lepton neutrinos outward. These effects carry over into the ejecta, where flavor conversions increase the low-$Y_{e}$ material by up to $300\%$ in the runs that have longer-lived NS remnants, translating into significantly larger relative r-process yields (up to a factor of 10). Note that the ejecta differences observed significantly exceed the reported $10\%$ to $30\%$ numerical uncertainties estimated from comparisons between M1 and Monte Carlo neutrino transport schemes Foucart et al. (2024). These differences are mainly driven by neutrino interactions in low-density regions, where neutrinos are less constrained by the surrounding matter and can oscillate more freely in flavor space.
All our models develop sparse regions unstable to FFIs and the patterns look qualitatively similar to those in Nagakura et al. (2025b). Notably, we find that the FFI-targeted flavor equilibration does not simply erase the instabilities. Unlike the more detailed, but local, calculations present in the literature, our global simulations allow us to capture non-local transport effects and dynamical changes in the neutrino distributions, which may contribute to the regeneration of instabilities. We emphasize that this conclusion should be validated with more sophisticated treatments of FFIs, particularly in accurately identifying their onset conditions and equilibrium states. While all simulations evolve toward flavor equilibrium points, the mixing models in general shift the flavor balance by enhancing heavy-lepton neutrino fractions at the expense of electron-type species. The density-11 mixing models, in particular, tend to produce more balanced fractions of $\nu_{e}$ and $\bar{\nu}_{e}$ compared to the FFI models, for short-lived remnant.
@@
As one of the first explorations of flavor conversion effects in neutron star mergers, our study emphasizes both the potential significance and the current uncertainties of such processes. We call for future efforts to build a more complete picture. A more comprehensive treatment of neutrino transport—including the effects of magnetic fields Tambe et al. (2025); Most et al. (2025), muonic interactions Pajkos and Most (2025b); Ng et al. (2024), inelastic neutrino-lepton scatterings Cheong et al. (2025); Chiesa et al. (2024), and pair processes—will be essential, as these missing ingredients can all have quantitative impacts. Equally important are theoretical advances in understanding the precise locations, timescales, and nonlinear outcomes of flavor instabilities, e.g. Richers et al. (2024); Liu et al. (2025b); Nagakura et al. (2025b); Laraib and Richers (2025); Liu et al. (2025c) (see also a review in Johns et al. (2025)), which are critical for accurately modeling flavor conversions in global merger simulations.

### v2

@@ III.3 $Y_{e}-\theta$ profiles
In Fig. 9 we show the average electron fractions for the ejecta as a function of the latitude $\theta$, for both the DD2 and SFHo runs. For the DD2 models, the ejecta average $Y_{e}$ increases almost monotonically with latitude. While for the SFHo runs, the average $Y_{e}$ of ejecta shows a small bump around $\theta\approx 20^{\circ}$, which is likely associated with shock heating. The main differences between the no-mixing and mixing runs with the DD2 EoS occur near the equatorial plane and at intermediate latitudes ($\theta<45^{\circ}$), with their average $Y_{e}$ differing by up to $0.1$. In the SFHo runs, the contrast between mixing and no-mixing models is likewise most pronounced in the near-equatorial ejecta, with little variation in the polar regions. The only exception is the SFHo-FFI model, which exhibits up to $\sim 0.05$ lower average $Y_{e}$ than the SFHo-NM model at higher latitudes ($\theta>45^{\circ}$). The differences in the $Y_{e}-\theta$ profiles between mixing and no-mixing models suggest that neutrino flavor conversions have larger impacts on ejecta near the equatorial plane and at intermediate latitudes, than ejecta in the polar regions. This is because the equatorial ejecta expand on timescales shorter than the weak equilibration timescale, whereas the polar ejecta expand more slowly and experience stronger neutrino irradiation, especially in the longer-lived remnant case. Therefore, the changes in neutrino luminosities and mean energies caused by flavor conversion have a more limited effect on the electron-fraction evolution of the polar ejecta.
@@ IV Conclusion
In agreement with earlier work Qiu et al. (2025); Li and Siegel (2021); Just et al. (2022), we find that mixing scenarios drive the matter toward more neutron-rich conditions compared to no-mixing models. The neutrino flavor transformations of electron (anti)neutrinos into heavy-lepton types in the inner disk regions boost the total neutrino luminosities at larger radii. While species-dependent hierarchies vary across the EoS and collapse outcomes, we generally find heavy-lepton luminosities amplified in the mixing models relative to the no-mixing case, and vice versa for the electron (anti)neutrinos. Flavor conversions raise the mean energies of electron (anti)neutrinos and lower those of heavy-lepton neutrinos, effectively shifting the decoupling surface of electron-flavor neutrinos inward and that of heavy-lepton neutrinos outward. These effects carry over into the ejecta, where flavor conversions increase the low-$Y_{e}$ material by up to $300\%$ in the runs that have longer-lived NS remnants, translating into significantly larger relative r-process yields (up to a factor of 10). Note that the ejecta differences observed significantly exceed the reported $10\%$ to $30\%$ numerical uncertainties estimated from comparisons between M1 and Monte Carlo neutrino transport schemes Foucart et al. (2024). While the enhancement in heavy-element production in the FFI model is on the order of 100%, which is consistent with the magnitude reported in Just et al. (2022). These differences are mainly driven by neutrino interactions in low-density regions, where neutrinos are less constrained by the surrounding matter and can oscillate more freely in flavor space.
All our models develop sparse regions unstable to FFIs and the patterns look qualitatively similar to those in Nagakura et al. (2025a). Notably, we find that the FFI-targeted flavor equilibration does not simply erase the instabilities. Unlike the more detailed, but local, calculations present in the literature, our global simulations allow us to capture non-local transport effects and dynamical changes in the neutrino distributions, which may contribute to the regeneration of instabilities. We emphasize that this conclusion should be validated with more sophisticated treatments of FFIs, particularly in accurately identifying their onset conditions and equilibrium states. While all simulations evolve toward flavor equilibrium points, the mixing models in general shift the flavor balance by enhancing heavy-lepton neutrino fractions at the expense of electron-type species. The density-11 mixing models, in particular, tend to produce more balanced fractions of $\nu_{e}$ and $\bar{\nu}_{e}$ compared to the FFI models, for short-lived remnant.
@@
As one of the first explorations of flavor conversion effects in neutron star mergers, our study emphasizes both the potential significance and the current uncertainties of such processes. We call for future efforts to build a more complete picture. A more comprehensive treatment of neutrino transport—including the effects of magnetic fields Tambe et al. (2025); Most et al. (2025), muonic interactions Pajkos and Most (2025a); Ng et al. (2024), inelastic neutrino-lepton scatterings Cheong et al. (2025); Chiesa et al. (2024), and pair processes—will be essential, as these missing ingredients can all have quantitative impacts. Equally important are theoretical advances in understanding the precise locations, timescales, and nonlinear outcomes of flavor instabilities, e.g. Richers et al. (2024); Liu et al. (2025c); Nagakura et al. (2025a); Laraib and Richers (2025); Liu et al. (2025a) (see also a review in Johns et al. (2025)), which are critical for accurately modeling flavor conversions in global merger simulations.